A variable name in a certain computer language must be either an alphabet or a alphabet followed by decimal digit.Total number of different variable names that can exist in that language is:


a) 280


b) 290


c)286


d)296


Plz post the full solution quickly.


ย 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear Priyansh


case1:  variable name has only alphabet

  total number of alphabet =26

case 2:  variable name has both alphabet and decimal

  total number = 26*10 =260


so total possibility = 260 + 26 = 286



Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin